How many molecules are there in 3 moles of aluminum chloride?


1
Expert's answer
2021-05-14T05:56:36-0400

Given n.o. of moles=3moles

To calculate n.o. of molecules =N

Use the Avogadro's constant =6.022×1023

Therefore:- N=3×6.022×1023

N=1.80066×1024Molecules